OpenStudy (uri):

Find the arrangements that can be made out of the letters of the words: 1)COMMERCE 2)STATISTICS

OpenStudy (anonymous):

what would it be if the letters were all distinct?

OpenStudy (uri):

I have no idea,I am new to this subject.

OpenStudy (anonymous):

oh ok counting principle

OpenStudy (anonymous):

suppose you had 8 distinct letters and you arrange them 8 choices for the first letter 7 for the second 6 for the third 5 for the 4th 4 for the 5th 3 for the 6th 2 for the 7th and finally 1 for the 8th by counting principle it is the product better known as \(8!\)

OpenStudy (uri):

8 letters two repeating so,(8.7.6.5.4.3.2.1)/2.1 then? @Mimi_x3

OpenStudy (anonymous):

not quite in commerce you have two C, two E, and two M

OpenStudy (uri):

Oh yeah..

OpenStudy (anonymous):

so \(8!\) over counts by a factor of \(8\)

OpenStudy (anonymous):

if you just wanted a formula it would be \[\frac{8!}{2!2!2!}\]

OpenStudy (anonymous):

but that is kind of silly since \(2!\) is just 2, so the denominator is 8, which cancels with the 8 up top giving you \(7!\)

OpenStudy (uri):

Oh 2! three times because 3 letters are repeating 2 times?

OpenStudy (anonymous):

right if for example you had googoo it would be \[\frac{6!}{2!4!}\]

OpenStudy (anonymous):

you should make sure you understand why that is

terenzreignz (terenzreignz):

there are THREE 2!'s in your evaluation, you only put 2 of them...

terenzreignz (terenzreignz):

No... it comes out different on google.

OpenStudy (uri):

OMG WAIT DUMB ME!! I LOVE YA TERENCE!

terenzreignz (terenzreignz):

:)

terenzreignz (terenzreignz):

Where'd you go wrong, though?

terenzreignz (terenzreignz):

ahh yeah I didn't see that too... d'oh XD

OpenStudy (uri):

Same i don't know where i am wrong,AM I EVEN WRONG?!! @nincompoop

terenzreignz (terenzreignz):

you put 2x2x2 as 6 Which is 2+2+2, not 2x2x2 :P

OpenStudy (uri):

Oh wait!!

OpenStudy (uri):

Okay so for COMMERCE it is 8!2!2! So that will be (8.7.6.5.4.3.2.1.)/(2.1)(2.1) So,40,320/8=5,040...Right?

terenzreignz (terenzreignz):

Yes. Much better :3

OpenStudy (uri):

Let's do *STATISTICS* now,I'll do you'll see? :3 Tell me where i do wrong.

terenzreignz (terenzreignz):

Me sleepy :3

OpenStudy (uri):

Sleep I'm doing...then you check later? :)

terenzreignz (terenzreignz):

yeah, sure... :) --------------------------- Terence out

OpenStudy (uri):

STATISTICS is 10!/3!3!2!